# DE-BERT: Distance-Enhanced Early Exiting for BERT based on Prototypical Networks

### v1

@@ IV-E2 Effect of DAR and Prototypical Networks
To investigate the effect of DAR and the prototypical networks, we plot the performance-efficiency trade-off curves of different models on a representative subset of GLUE, as shown in Fig. 5. All models adopt the proposed hybrid exiting strategy during inference. Implementation details can be found in Appendix B. We can see that removing either DAR or the prototypical networks causes significant performance degradation. The experimental results confirm the effectiveness of the prototypical networks and DAR in our framework, as they are indispensable for learning high-quality class prototype representations and metric space. This, in turn, plays a crucial role in accurately estimating the correctness of early predictions and making reliable exiting decisions. For DAR, the visualization results of sample representations demonstrate that our proposed DAR can effectively promote the learning of high-quality metric space and class prototype representations, which can be found in Section V-E. For the prototypical network, further correlation analysis experiments in Section V-F reveal that the introduction of prototypical networks effectively prevents the homogenization of entropy and distance ratio, facilitating their mutual correction.
V In-depth Analysis
In this section, we conduct an in-depth analysis of the parameter sensitivity, interpretability, storage and computational costs, generality, and the contribution of each component of our DE-BERT model. Most experiments are conducted on a representative subset of GLUE, including MNLI, QNLI, QQP, and SST-2 tasks, to guarantee consistent and reliable results.
V-A Impact of and
In this section, we conduct experiments on two representative tasks, i.e., SST-2 and QNLI, to investigate the impact of two parameters: in Equation (11) and in Equation (5).
V-A1 Impact of
The parameter in Equation (11) balances the contribution of entropy (local information) and distance ratio (global information) in the exit decision-making process. A higher value of indicates a stronger emphasis on the distance ratio (global information). We train the models as illustrated in Section III-B with a fixed value: 0.1 for SST-2, and 0.001 for QNLI, and then evaluate the performance-efficiency trade-off under different values of during inference. Fig. 6 shows the impact of on task performance under different speed-up ratios. Note that the experimental results in Fig. 6 include the performance improvements brought by DAR, and we only focus on the impact of on the model’s acceleration performance during inference.
Overall, we observe that, as the value of increases, the task performance exhibits a consistent pattern of initially increasing and then decreasing under different speed-up ratios. This indicates that the incorporation of the distance ratio can consistently improve the acceleration performance of the model under various speed-up ratios. It also suggests an optimal trade-off between entropy and distance ratio. Additionally, it is noticeable that as the speed-up ratio increases, the optimal value of tends to increase, and compared to entropy-based exiting strategies (), the introduction of distance ratio yields more significant performance improvements. We believe that this observation can be attributed to the following reasons. Under high speed-up ratios, samples tend to exit at shallow layers (see Appendix C-A). The limited capability of shallow classifiers amplifies the gap between the confidence level reflected by entropy and the correctness of early predictions. This makes it crucial to introduce the distance ratio (global information) to improve the estimation of prediction correctness, thus enhancing the reliability of exiting decisions. Consequently, the performance gains brought by distance ratio tend to be more significant. This explanation is subsequently demonstrated through experiments, as detailed in Section V-B. Finally, for parameter selection, the optimal value differs across tasks and speed-up ratios, as shown in Fig. 6. Therefore, it is recommended to conduct a parameter search within the range between 0 and 10 according to specific scenarios.
V-A2 Impact of
The parameter in Equation (5) balances the cross-entropy loss and DAR during training. A higher value of indicates a stronger emphasis on optimizing DAR. We train the models as illustrated in Section III-B, with values of {0.0001, 0.001, 0.01, 0.1}. During inference, the proposed hybrid exiting strategy is employed with a fixed value of 1.0 for SST-2 and 2.0 for QNLI. Fig. 7 shows the impact of on task performance under different speed-up ratios. It can be observed that the optimal value of differs across tasks and speed-up ratios, but the performance changes caused by different values are always less than 1, indicating that the acceleration performance is not significantly affected by selection. Moreover, an value between 0.0001 and 0.1 can always lead to satisfactory performance on different tasks.
V-B Accuracy of Prediction Correctness Estimation
Overall, we observe that the value consistently outperforms the entropy in estimating the correctness of early predictions across different layers and tasks. This demonstrates that distance-based global information can complement entropy-based local information, offering a more accurate estimation of prediction correctness. Additionally, it is noticeable that as the number of layers decreases, the accuracy of prediction correctness estimation based on both entropy and values tends to decline, and the introduction of distance ratio yields more significant improvements in estimating prediction correctness. We believe that this observation can be attributed to the following reasons. The limited representation capability of internal classifiers in shallow layers leads to insufficient local information based on entropy, exacerbating the difficulty of estimating prediction correctness. Consequently, compared to deeper layers, introducing distance ratio at shallow layers is particularly crucial for enhancing the accuracy of prediction correctness estimation. This, in turn, contributes to more reliable exiting decisions and better trade-offs between model performance and efficiency. This further explains our observations in Fig. 4 and Fig. 6, that as the speed-up ratio increases, the introduction of distance ratio tends to yield greater performance improvements.
V-C Storage and Computational Costs
V-C1 Parameter Volumes and Training Time Costs
Table IV presents the parameter volumes and training time costs of our DE-BERT model and those of DeeBERT [20] for comparison. As shown in Table IV, compared to the classic entropy-based early exiting model DeeBERT, our DE-BERT model requires only an additional 5.9 of parameters due to the incorporation of prototypical networks. Notably, the number of additional parameters is irrelevant to the number of classes, which is attributed that the introduced prototypical networks are -agnostic. Furthermore, despite additional parameters, the training time costs are not significantly increased and are even reduced in some tasks such as MRPC and QQP. This is attributed that the additional supervision provided by DAR accelerates the model convergence during training. Therefore, our method can achieve a better trade-off between model performance and efficiency during inference with minimal additional parameters and comparable training time costs, compared to the classic entropy-based early exiting model.
| Model | Params | Training Time (min) | |||||
| MRPC | SST-2 | QQP | MNLI | ||||
| DeeBERT | 109.50M | 109.51M | 1.3 | 34 | 304 | 163 | |
| DE-BERT (ours) | 6.50M | 6.50M | 0.95 | 44 | 293 | 165 | |
V-C2 Computational Complexity
Table V shows each module’s computational complexity within the DE-BERT model. We can observe that, due to the substantial self-attention operations, the computational complexity of each encoder block is 1813.5M, comprising the major part of the model’s computational costs. In contrast, the computational complexity of each prototypical network is only 1.2M, leading to a 0.066 increase in computational costs per layer. Consequently, the incorporation of prototypical networks can enhance the reliability of exiting decisions with negligible additional computational costs arising from distance calculation. Moreover, it can be observed that the additional computational costs provided by prototypical networks are irrelevant to the number of classes , indicating that the prototypical networks are -agnostic. Furthermore, as shown in Table V, the computational complexity of an encoder block significantly surpasses that of other modules. Therefore, we can draw a conclusion that the model’s computational costs during inference are approximately proportional to the number of executed layers, confirming the rationality of the speed measurement illustrated in Equation (12). Further exploration of this point is provided in Appendix C-B.
| Module | FLOPs | |
| Embedding | 786.4K | 786.4K |
| Encoder | 1813.5M | 1813.5M |
| Pooler | 1.2M | 1.2M |
| Prediction Head | 3.1K | 4.6K |
| Prototypical Network | 1.2M | 1.2M |
| Internal Classifier | 3.1K | 4.6K |
V-D Generality on Different PLMs
To investigate the generality of our method on different PLMs, we conduct experiments on a representative subset of GLUE with RoBERTa [3] as the backbone model. RoBERTa is an optimized PLM based on the BERT architecture. Compared to BERT, RoBERTa utilizes more training data, larger batch sizes, and better training strategies, leading to improved performance. Table VI shows the classification performance of each early exiting method under the speed-up ratio 3.0. We can observe that our method outperforms all competing early exiting methods in general, demonstrating its potential as a generic method for accelerating the inference of different types of PLMs.
@@ V-D Generality on Different PLMs
| GPFEE | 81.4 | 89.2 | 71.9 | 93.5 | 84.0 |
| DE-RoBERTa (ours) | 83.1 | 89.3 | 71.6 | 93.7 | 84.4 |
V-E Visualization of Sample Representations
To verify whether DAR facilitates the learning of metric space and class prototype representations, we use the t-SNE projection [54] to visualize the sample representations generated by the prototypical networks of the DE-BERT model. For comparison, we train a variant (denoted by DE-BERT w/o DAR) of the DE-BERT model using only cross-entropy loss and visualize the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers). Fig. 9 shows the visualization results on layers 2 and 5 of the two models on the MNLI development set. We can observe that, on both layers 2 and 5, DAR encourages clearer classification boundaries between different classes. This indicates that DAR can facilitate the learning of high-quality metric space and class prototype representations, which is crucial for accurately estimating the prediction correctness and making reliable exiting decisions, as shown in Fig. 5.
V-F Additional Effect of Prototypical Networks
As mentioned in Section III-A, the main purpose of introducing the prototypical networks is to learn the class prototype representations and the associated metric space. Additionally, we believe that the introduction of prototypical networks can help avoid the homogeneity issue between entropy and distance ratio since they are computed in different representation spaces. This enables them to mutually correct each other. To verify our claim, we analyze the Spearman correlation coefficient between the entropy and distance ratio of our DE-BERT models with and without prototypical networks, respectively. Concretely, we train a variant (denoted by DE-BERT w/o PN) of our DE-BERT model as illustrated in Section III-B. The computation of DAR terms and the updating of class prototype representations during training, as well as the calculation of distance ratios during inference, are all based on the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers). We then analyze the Spearman correlation coefficient between the entropy and distance ratio of this variant and the full DE-BERT model. Table VII illustrates the analysis results on a representative subset of GLUE based on the model outputs at the second layer. The results indicate a significant increase in correlation between entropy and distance ratio when the prototypical networks are removed. These results are consistent with the results shown in Fig. 5 that removing the prototypical networks impairs the acceleration performance of the model, which further confirms that the prototypical networks can effectively avoid the homogeneity between entropy and distance ratio, thus enhancing their integration.
| Model | QQP | MNLI | SST-2 | QNLI |
| DE-BERT | 0.207 | 0.018 | 0.145 | 0.066 |
| DE-BERT w/o PN | 0.569 | 0.389 | 0.479 | 0.422 |
V-G Experimental Results of Different DAR Forms
Table VIII shows the performance comparison of different DAR forms on the GLUE development sets under various speed-up ratios. Overall, the experimental results show that different forms of DAR yield comparable results, which confirms the generality of our framework across various metric learning methods. Specifically, as the speed-up ratio increases, Center Loss exhibits a slight advantage over the other two DAR forms in terms of performance degradation, yielding a better trade-off between model performance and efficiency, although the performance gaps are not significant. We speculate that this is caused by the conflicting optimization of Alienation Loss and cross-entropy loss during the training phase. Concretely, both Alienation Loss and cross-entropy loss have the effect of enlarging the difference between intra-class and inter-class distances, but their different implementation mechanisms lead to conflicting optimization during training. As a result, this affects the learning of metric space and class prototype representations, which ultimately compromises the acceleration performance of the model. In contrast, Center Loss, which aims to reduce intra-class distances, is better suited to cross-entropy loss. This facilitates the learning of high-quality metric space and class prototype representations, thus improving the acceleration performance of the model. Therefore, considering the above analysis and the computational efficiency, we primarily use Center Loss in DAR.
| Method | MNLI | MRPC | QNLI | QQP | RTE | SST-2 | AVG | |
| (393K) | (3.7K) | (105K) | (364K) | (2.5K) | (67K) | |||
| BERT- | 83.5 (1.00) | 87.5 (1.00) | 91.2 (1.00) | 89.8 (1.00) | 71.1 (1.00) | 91.5 (1.00) | 85.8 | |
| 1.33 | DE-BERT | 83.4 (1.41) | 88.0 (1.36) | 90.6 (1.39) | 91.1 (1.38) | 69.5 (1.45) | 91.8 (1.34) | 85.7 |
| DE-BERT-v2 | 83.2 (1.39) | 87.1 (1.38) | 91.1 (1.38) | 91.1 (1.37) | 69.8 (1.44) | 91.9 (1.33) | 85.7 | |
| DE-BERT-v3 | 83.2 (1.41) | 87.3 (1.34) | 91.1 (1.39) | 91.2 (1.39) | 69.7 (1.42) | 91.8 (1.32) | 85.8 | |
| 2 | DE-BERT | 82.7 (2.13) | 85.3 (2.13) | 90.0 (2.18) | 91.1 (2.06) | 69.6 (1.99) | 91.3 (1.96) | 85.0 |
| DE-BERT-v2 | 82.7 (2.14) | 84.6 (2.15) | 90.3 (2.17) | 91.1 (1.98) | 69.7 (2.07) | 91.5 (1.99) | 84.9 | |
| DE-BERT-v3 | 82.6 (2.13) | 85.1 (2.13) | 90.1 (2.18) | 91.2 (1.99) | 69.7 (1.99) | 91.4 (2.02) | 85.0 | |
| 4 | DE-BERT | 73.8 (4.37) | 79.2 (4.35) | 81.1 (4.02) | 88.0 (5.52) | 60.4 (4.38) | 89.6 (4.36) | 78.7 |
| DE-BERT-v2 | 74.0 (4.38) | 81.6 (4.36) | 80.5 (4.05) | 87.4 (5.50) | 58.5 (4.37) | 89.3 (4.35) | 78.6 | |
| DE-BERT-v3 | 74.0 (4.36) | 80.3 (4.36) | 80.7 (3.97) | 87.5 (5.54) | 58.8 (4.35) | 89.4 (4.37) | 78.5 |
VI Discussion
VI-A Discussion of Experimental Results
In general, the experimental results demonstrate our hypothesis that the performance bottleneck of existing early exiting methods primarily stems from their exclusive reliance on local information from an individual test sample to make exiting decisions, while overlooking the global information provided by the sample population. We believe that the latent class information of a test sample can be reflected by its distances from class prototypes, which can effectively complement the entropy-based local information and improve the accuracy of prediction correctness estimation from a global perspective. In this way, the proposed DE-BERT approximates the behavior of the oracle, thus achieving a better trade-off between model performance and efficiency. Furthermore, the limited representation capability of shallow internal classifiers leads to insufficient local information based on entropy, which emphasizes the importance of introducing global information. This explains the superiority of our hybrid exiting strategy in high acceleration scenarios.
VI-B Limitations and Future Work
For further research in the near future, we comprehensively discuss the limitations of this work from three perspectives. Firstly, our DE-BERT framework caters to classification tasks only since DE-BERT incorporates the distance ratio that requires the prototype representations of all classes, which is unavailable for other non-classification tasks. One of our future studies is to extend our DE-BERT framework to more sophisticated regression tasks by training a neural network to map continuous regression values to their corresponding prototype representations. Secondly, our framework DE-BERT improves the entropy-based early exiting by transferring knowledge from the training set to the test set via class prototypes, assuming that the distribution is identical between the training and test data, i.e., in-distribution generalization. The out-of-distribution case brings extreme challenges to DE-BERT, which is worth further research and exploration. Finally, different from our DE-BERT which focuses on improving the exiting strategies, some research, such as CascadeBERT [50], GPFEE [26], and LeeBERT [29], primarily address different issues of early exiting, including training schemes, model architecture, and model calibration. Exploring the combination of our DE-BERT with the aforementioned methods would be an interesting endeavor, as it contributes to further enhancing the acceleration performance of early exiting models by investigating the potential orthogonality and complementarity between different techniques.
VII Conclusion
@@
In this article, we point out that the performance limitation of existing early exiting methods primarily lies in their excessive focus on local information from an individual test sample while neglecting the global information offered by the sample population, which compromises the estimation of prediction correctness, leading to erroneous exiting decisions. To remedy this, we propose the DE-BERT framework, which leverages prototypical networks to provide distance-based global information, improving the estimation of prediction correctness for reliable early exiting. Our framework is both intuitive and interpretable. Extensive experiments on the GLUE benchmark demonstrate the superiority of our framework under various speed-up ratios with minimal additional storage and computational costs. Further analysis also validates our framework’s interpretability, insensitivity to parameters, and generality on different PLMs.
@@ A-1 Backbone
BERT-base [2] consists of 12 encoders with a single-layer classifier attached to the last encoder. With 110M parameters, a hidden size of 768, and 12 attention heads per layer, it is first pre-trained on the Wiki corpus and then fine-tuned on specific tasks.
@@ A-3 Knowledge Distillation
Knowledge distillation methods compress models by transferring knowledge from a larger teacher model to a more compact student model. DistilBERT [46] is distilled during the pre-training phase. PD-BERT [47] is first pre-trained on a corpus and then distilled during the fine-tuning stage. BERT-PKD [48] is distilled using both the final predictions and intermediate-layer outputs provided by the teacher model during the fine-tuning stage. BERT-of-Theseus [49] is distilled using a progressive module replacement strategy during the fine-tuning stage. All of the compressed models mentioned above consist of 6 encoders, resulting in a speed-up ratio of 2.0.
@@ A-4 Early Exiting
Regarding the inference stage, DeeBERT [20] and GPFEE [26] employ confidence-based exiting strategies, which use the entropy of early predictions to reflect the confidence level of internal classifiers. The inference process is terminated once the entropy value falls below a predefined threshold. PABEE [21] and LeeBERT [29] both employ a patience-based exiting strategy, which utilizes cross-layer consistency as the exiting indicator. The exiting criterion is met when there are enough (i.e., the patience threshold) consecutive internal classifiers that agree with each other. BERxiT [22] incorporates the LTE module, a one-layer fully-connected network that learns to score the correctness of early predictions. Early exit is triggered once the correctness score surpasses the threshold. PCEE-BERT [23] employs a hybrid exiting strategy that combines confidence-based and patience-based strategies. It uses entropy as the confidence measure and terminates the inference if there are enough numbers (i.e., the patience threshold) of consecutive internal classifiers being confident for their predictions.
@@
Regarding the architectures of internal classifiers, GPFEE [26] enhances the internal classifiers by integrating both past and future states as its inputs to facilitate high-quality early predictions, while the remaining models utilize a single-layer fully-connected network as the internal classifier.
@@
Regarding the training phase, DeeBERT [20], PABEE [21], and PCEE-BERT [23] are trained by minimizing the cross-entropy losses for all classifiers. For the remaining models, despite the cross-entropy loss for internal classifiers, additional training objectives are needed. BERxiT [22] introduces the MSEs for all LTE modules to encourage high-quality scoring results for prediction correctness. GPFEE [26] leverages the imitation loss for generating approximations of future states based on all available past states. LeeBERT [29] leverages a cross-layer distillation objective with learned weights to encourage mutual learning among internal classifiers, thus improving the training process.
The above information is summarized in Table IX.
Appendix B Implementation Details for Ablation Studies
To verify the effectiveness of the prototypical networks (PNs) and DAR, we train two variants of our DE-BERT model and employ the proposed hybrid exiting strategy for these models as well as the full DE-BERT model during inference. We then compare the performance-efficiency trade-offs of different models.
B-1 DE-BERT w/o PN
To demonstrate the effectiveness of the prototypical networks, we design and train a variant of the DE-BERT model (denoted by DE-BERT w/o PN) as illustrated in Section III-B in the main paper. During inference, we use the proposed hybrid exiting strategy. Note that the computation of DAR terms and the updating of class prototype representations during training, as well as the calculation of distance ratios during inference, are all based on the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers).
B-2 DE-BERT w/o PN w/o DAR
To further demonstrate the effectiveness of DAR, we also train another variant of the DE-BERT model (denoted by DE-BERT w/o PN w/o DAR) only using the cross-entropy loss as the training objective. To ensure the feasibility of the proposed hybrid exiting strategy during inference, the class prototype representations are still updated during training. Note that the updating of class prototype representations during training, as well as the calculation of distance ratios during inference, are all based on the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers).
@@ B-2 DE-BERT w/o PN w/o DAR
| LeeBERT | Patience-based | Single-layer FC | , |

### v2

@@ IV-E2 Effect of DAR and Prototypical Networks
To investigate the effect of DAR and the prototypical networks, we plot the performance-efficiency trade-off curves of different models on a representative subset of GLUE, as shown in Fig. 5. All models adopt the proposed hybrid exiting strategy during inference. Implementation details can be found in Appendix B. We observe that removing either DAR or the prototypical networks causes significant performance degradation, demonstrating their effectiveness in learning high-quality class prototypes and metric spaces. This is crucial for accurately estimating the prediction correctness and making reliable exiting decisions. We further visualize the effect of DAR in learning high-quality metric space and class prototypes (see Appendix C-H). Additionally, we reveal that the prototypical networks can prevent the homogenization of entropy and distance ratios, facilitating their mutual correction (see Appendix C-I).
IV-F Generality on Different PLMs
To investigate the generality of our method on different PLMs, we conduct experiments on a representative subset of GLUE with RoBERTa [abs-1907-11692] as the backbone model, as shown in Table IV. RoBERTa is an optimized variant of BERT that utilizes more training data, larger batch sizes, and better training techniques. The experimental results indicate that our method generally outperforms all early exiting baselines, and the one-tailed t-tests further verify the statistical significance of our method’s improvements over the competitive baseline GPFEE [globalpast] in most tasks. These findings demonstrate the generality of our framework across different PLMs.
@@ IV-F Generality on Different PLMs
| GPFEE† | 81.4 | 89.2 | 71.9 | 93.5 | 84.0 |
| DE3-RoBERTa (ours) | 83.1∗ | 89.3∗ | 71.6 | 93.7∗ | 84.4∗ |
V Conclusion
@@
In this article, we point out that the performance limitation of existing early exiting methods primarily lies in their excessive focus on local information from individual test samples while neglecting the global information offered by the sample population, which compromises the estimation of prediction correctness, leading to erroneous exiting decisions. To remedy this, we propose the DE3-BERT framework, which leverages prototypical networks to provide distance-based global information, improving the estimation of prediction correctness for reliable early exiting. Our framework is both intuitive and interpretable. Extensive experiments on the GLUE benchmark demonstrate the superiority of our framework under various speed-up ratios with minimal additional storage and computational costs. Further analysis validates the effectiveness and interpretability of each component in our framework, as well as its robustness to hyperparameter selection and its strong generality across out-of-distribution (OOD) data, different languages, and various backbone models. We also provide a detailed discussion of the experimental results, limitations, scalability, and real-world applications and implications of our framework (see Appendix E).
@@ A-1 Backbone
BERT-base [bert] consists of 12 encoders with a single-layer classifier attached to the last encoder. With 110M parameters, a hidden size of 768, and 12 attention heads per layer, it is first pre-trained on the Wiki corpus and then fine-tuned on specific tasks.
@@ A-3 Knowledge Distillation
Knowledge distillation methods compress models by transferring knowledge from a larger teacher model to a more compact student model. DistilBERT [distilbert] is distilled during the pre-training phase. PD-BERT [wellread] is first pre-trained on a corpus and then distilled during the fine-tuning stage. BERT-PKD [PKD] is distilled using both the final predictions and intermediate-layer outputs provided by the teacher model during the fine-tuning stage. BERT-of-Theseus [theseus] is distilled using a progressive module replacement strategy during the fine-tuning stage. All of the compressed models mentioned above consist of 6 encoders, resulting in a speed-up ratio of 2.0.
@@ A-4 Early Exiting
Regarding the inference stage, DeeBERT [Deebert], GPFEE [globalpast], and COSEE [COSEE] employ confidence-based exiting strategies, which use the entropy or energy value of early predictions to reflect the confidence level of internal classifiers. The inference process is terminated once the entropy or energy value falls below a predefined threshold. PABEE [Pabee] and LeeBERT [zhu2021leebert] both employ a patience-based exiting strategy, which utilizes cross-layer consistency as the exiting indicator. The exiting criterion is met when there are enough (i.e., meeting the patience threshold) consecutive internal classifiers that agree with each other. BERxiT [Berxit] incorporates a learning-to-exit network to score the correctness of early predictions. Early exit is triggered once the correctness score surpasses the threshold. PCEE-BERT [PCEE] employs a hybrid exiting strategy that combines confidence-based and patience-based strategies. It uses entropy as the confidence measure and terminates the inference if there are enough (i.e., meeting the patience threshold) consecutive internal classifiers that are confident in their predictions.
@@
Regarding the architecture of internal classifiers, GPFEE [globalpast] enhances the internal classifiers by integrating both past and future states as its inputs to facilitate high-quality early predictions, while the remaining models utilize a single-layer fully-connected network as the internal classifier.
@@
Regarding the training phase, DeeBERT [Deebert], PABEE [Pabee], and PCEE-BERT [PCEE] are trained by minimizing the cross-entropy losses for all classifiers. For the remaining models, despite the cross-entropy loss for internal classifiers, additional training objectives are needed. BERxiT [Berxit] introduces mean squared error (MSE) for the learning-to-exit network to encourage high-quality scoring results for prediction correctness. GPFEE [globalpast] leverages the imitation loss for generating approximations of future states based on all available past states. LeeBERT [zhu2021leebert] leverages a cross-layer distillation objective with learned weights to encourage mutual learning among internal classifiers. COSEE [COSEE] introduces the online signal calibration loss to produce highly discriminative exiting indicators, facilitating the reliability of exiting decisions.
The above information is summarized in Table V.
@@
| LeeBERT | Patience-based | Single-layer FC | , | |
| COSEE | Confidence-based | Single-layer FC | , |
Appendix B Implementation Details for Ablation Studies
To verify the effectiveness of the prototypical networks (PNs) and DAR, we train two variants of our DE3-BERT model and employ the proposed hybrid exiting strategy for these models as well as the full DE3-BERT model during inference. We then compare the performance-efficiency trade-offs of different models.
B-1 DE3-BERT w/o PN
To demonstrate the effectiveness of the prototypical networks, we design and train a variant of the DE3-BERT model (denoted by DE3-BERT w/o PN) as illustrated in Section III-B in the main paper. During inference, we use the proposed hybrid exiting strategy. Note that the computation of DAR terms and the updating of class prototype representations during training, as well as the calculation of distance ratios during inference, are all based on the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers).
B-2 DE3-BERT w/o PN w/o DAR
To further demonstrate the effectiveness of DAR, we also train another variant of the DE3-BERT model (denoted by DE3-BERT w/o PN w/o DAR) only using the cross-entropy loss as the training objective. To ensure the feasibility of the proposed hybrid exiting strategy during inference, the class prototype representations are still updated during training. Note that the updating of class prototype representations during training, as well as the calculation of distance ratios during inference, are all based on the sample representations obtained by pooling the hidden states of encoder layers (i.e., inputs to internal classifiers).